6. What amount must be invested to accumulate to a value of S11 900 in 180 days if the interest rate is 7%? a) S11 502.91 b) S11 489.20 c) S11 487.58 d) S11 603.48 7. You have invested $2500 in a 30-day savings certificate paying 4.15% per annum. Find the amount of the investment at the end of 30 days. a) S2551.88 b) S2508.53 c) S2491.50 d) S2603.75 8. Consider a payment of S600 due in six months from today. Find the equivalent value at two months if the interest rate per annum is 5.5 % a) $600.00 b) $589.20 c) S594.55 d) S583.94
Answer #1

(6) Let the required investment be $ M

Target Future Value = $ 11900, Tenure of Investment = 180 days and Required Interest Rate = 7 %

Therefore, M x [1+(0.07) x (180/365)] = 11900

M = 11900 / 1.034521 ~ $ 11502.91

Hence, the correct option is (a)

(7) Initial Investment = $ 2500, Interest Rate = 4.15 % per annum and Investment Tenure = 30 days

Future Value of Investment = 2500 x [1+(0.0415) x (30/365)] = $ 2508.527 ~ $ 2508.53

Hence, the correct option is (b)

(8) Interest Rate = 5.5 % per annum, Tenure = 6 - 2 = 4 months and FUture Value = $ 600

Present value at 2 months = 600 / [1+(0.055) x (2/12)] = $ 589.198 ~ $ 589.2

Hence, the correct option is (b)